# On the controller form for linear hyperbolic MIMO systems with dynamic boundary conditions

### v1

@@ IV-D4 Algorithm
The matrix $\bar{}\mathbf{H}(\sigma,\partial_{t})=\mathbf{L}(\sigma,\partial_{t})\mathbf{H}(\sigma,\partial_{t})$ obtained by LABEL:alg:shift_red satisfies the requirements (24).
(25).
@@
Starting in the second row, the degrees in each column are corrected from left to right, where each of these correction steps will introduce new undesired shifts to its left in the rows below its diagonal element. Therefore, the procedure then continues to the next row, until no more new shifts are introduced as the last row is reached. ∎
IV-E Hyperbolic MIMO controller form
Applying LABEL:alg:shift_red to $\mathbf{H}(\sigma,\partial_{t})$, we obtain
$\mathbf{L}^{-1}(\sigma,\partial_{t})\bm{u}(\partial_{t})=\bar{}\mathbf{H}(\sigma,\partial_{t})\bm{y}(\partial_{t})$ (31)
with $\hat{\mathbf{H}}=\operatorname{lccm}_{s,\sigma}\bar{}\mathbf{H}(\sigma,\partial_{t})$ having full rank. Thus, by means of the separation $\bar{}\mathbf{H}(\sigma,\partial_{t})=\hat{}\mathbf{H}\mathbf{K}(\sigma,\partial_{t})+\tilde{}\mathbf{H}(\sigma,\partial_{t})$ where $\mathbf{K}(\sigma,\partial_{t})=\operatorname{diag}(\partial_{t}^{\nu_{1}}\sigma^{\hat{\tau}_{1}},\dotsc,\partial_{t}^{\nu_{m}}\sigma^{\hat{\tau}_{m}})$ with the highest derivative and prediction orders $\nu_{i}=\deg_{\partial_{t}}\bar{h}_{i,i}(\sigma,\partial_{t})$ and $\hat{\tau}_{i}=\deg^{+}_{\sigma}\bar{h}_{i,i}(\sigma,\partial_{t})$ we obtain the explicit DDE system
$\mathbf{K}(\sigma,\partial_{t})\bm{y}(\partial_{t})=\hat{\mathbf{H}}^{-1}\!\left(\mathbf{L}^{-1}(\sigma,\partial_{t})\bm{u}(\partial_{t})\!-\tilde{\mathbf{H}}(\sigma,\partial_{t})\bm{y}(\partial_{t})\right)$ (32)
@@ IV-E Hyperbolic MIMO controller form
$y_{i}^{(\nu_{i})}(t+\hat{\tau}_{i})=-\mathcal{A}_{i}(t)*\bm{y}(t)+\mathcal{B}_{i}(t)*\bm{u}(t)$ (33)
for $i=\{1,\dotsc,m\}$ in the analytic setting. Herein, the elements of $\mathcal{A}_{i},\mathcal{B}_{i}$ by construction have compact support, which is why the convolutions will yield definite integrals whose intervals depend on the degrees of $\tilde{\mathbf{H}}(\sigma,\partial_{t})$ and $\mathbf{L}^{-1}(\sigma,\partial_{t})$ w.r.t $\sigma$. Thus, (33) states a well posed initial value problem if the components $y_{i}(t+\tau_{i})$ of the flat output are known for $\tau_{i}\in[\check{\tau}_{i},\hat{\tau}_{i}]$ with $\check{\tau}_{i}=\deg^{-}_{\sigma}\bar{h}_{i,i}(\sigma,\partial_{t})$ and $\bm{u}$ is known on $\mathbb{R}$.
@@
$\eta^{\mathrm{O}}_{i,j}(t)$ $\coloneq y_{i}^{(j-1)}(t-\check{\tau}_{i}),\;j\in\{1,\dotsc,\nu_{i}\}$ (34a)
$\eta^{\mathrm{P}}_{i}(\tau,t)$ $\coloneq y_{i}^{(\nu_{i})}(t+\tau),\;\tau\in[\check{\tau}_{i},\hat{\tau}_{i}]$ (34b)
for $i\in\{1,\dotsc,m\}$ which is equivalent to the flat output on these intervals. Substituting [^5] Some expressions in $\bm{y}(t)$ may remain as e.g. the values $y_{i}^{(d)}(t+\tau_{i})$ are not part of the HCF state $\bm{\eta}(\cdot,t)$ for $p<\nu_{i}$, cf. (12). However, these expressions can be easily eliminated by means shown e.g. in [6, Lemma 5]. the definitions (34) in (33) then allows to identify the functions $a^{\star}_{i,j}(t),b^{\star}_{i,j}(t)$ as well as weights $\bm{a}_{i}$ in (12), yielding the HCF of system (15).
@@
The proof of Lemma 2 yields $\deg_{\partial_{t}}q(\sigma,\partial_{t})<0$ due to the structure of $d(\partial_{t})$. Then by Lemma 3, $q^{\star}(\sigma,\partial_{t})$ has non-positive degree as do all entries in $\mathbf{L}_{i}(\sigma,\partial_{t})$ from LABEL:alg:shift_red. Now, as the inverse of the triangular $\mathbf{L}(\sigma,\partial_{t})$ is computed by forward substitution, all $\bar{q}_{i,j}(\sigma,\partial_{t})$ therein are linear combinations of the $q_{i,j}^{\star}(\sigma,\partial_{t})$, hence, $\deg_{\partial_{t}}\bar{q}_{i,j}(\sigma,\partial_{t})\leq 0$. Therefore, the convolution with $\mathcal{B}_{i}$ yields no derivatives. ∎
@@ V Reduction of the benchmark system
Consider again the system from Section II-C with $\alpha=\pi$, $\beta=10$ and $m=1$ substituted into (10) yielding
$\mathbf{H}(\sigma,\partial_{t})=\frac{1}{4}\begin{pmatrix}a_{1}(\partial_{t})\sigma^{\pi}+a_{2}(\partial_{t})\sigma^{-\pi}&-\sigma^{\pi}-\sigma^{-\pi}\\ b_{1}(\partial_{t})\sigma^{10}+b_{2}(\partial_{t})\sigma^{-10}&\sigma^{10}+\sigma^{-10}\\ \end{pmatrix}$
with $a_{1}(\partial_{t})=b_{1}(\partial_{t})=\partial_{t}^{2}+2\partial_{t}$ and $a_{2}(\partial_{t})=b_{2}(\partial_{t})=\partial_{t}^{2}-2\partial_{t}$ for which $\operatorname{lccm}_{\sigma,\partial_{t}}\mathbf{H}(\sigma,\partial_{t})=\operatorname{lccm}_{\partial_{t},\sigma}\mathbf{H}(\sigma,\partial_{t})$ is of rank $1$ where $2$ would be needed.
Now, regarding Section IV-D1, the rows are already correctly ordered since $\pi<10$ and also the columns do not need to be reordered since the coefficients in each column are of the same degree w.r.t. $\partial_{t}$. Continuing with Section IV-D2, the conditions (25) fail for $h_{2,1}(\sigma,\partial_{t})$ as e.g. $\deg_{\sigma}h_{2,1}(\sigma,\partial_{t})=20>\deg_{\sigma}h_{1,1}(\sigma,\partial_{t})=2\pi$. Therefore, by means of Lemma 2 the quotient
$\begin{aligned} q_{1,2}&(\sigma,\partial_{t})\!=\!\sigma^{10-\pi}\!+\!\frac{2-\partial_{t}}{\partial_{t}+2}\sigma^{10-3\pi}\!-\!\frac{\partial_{t}+2}{\partial_{t}-2}\sigma^{-10+3\pi}\!+\!\sigma^{-10+\pi}\end{aligned}$
is obtained, where one iteration was needed for each case in (28). Next, Lemma 3 yields the correction term $p(\partial_{t})=\frac{16e^{-20+6\pi}}{\left(\partial_{t}^{2}-4\right)}$ and thus $q_{2,1}^{\star}(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{w}}$. Using the latter in (26) then yields the reduced element
$\begin{aligned} \bar{h}_{1,2}(\sigma,\partial_{t})=\!-\frac{16\partial_{t}}{\left(\partial_{t}-2\right)e^{20-6\pi}}\sigma^{\pi}\!+\frac{\partial_{t}\left(\partial_{t}^{2}+4\partial_{t}+4\right)}{\partial_{t}-2}\sigma^{-10+4\pi}\\ \!+\frac{\partial_{t}\left(\partial_{t}^{2}-4\partial_{t}+4\right)}{\partial_{t}+2}\sigma^{10-4\pi}\!-\frac{16\partial_{t}}{\left(\partial_{t}+2\right)e^{20-6\pi}}\sigma^{-\pi}.\end{aligned}$
As the requirements for the second column are already satisfied, this concludes the reduction and we obtain $\hat{}\mathbf{H}=\operatorname{lccm}_{\partial_{t},\sigma}\bar{}\mathbf{H}(\sigma,\partial_{t})=\operatorname{diag}(1,-2)$.
Python code showing the complete computation of the HCF will be made available at a later time.
@@ VI Concluding remarks
This contribution establishes a HCF for MIMO systems with dynamic boundaries and proposes a flatness based algorithm to compute said form for a specific class of systems. While the algebraic approach taken to obtain this result is limited to linear systems, it enables a systematic procedure which lends well to an implementation using computer algebra. However, the insight gained from an analytic interpretation of the algebraic procedure allows for the technique to be extended to nonlinear boundary systems, which will be discussed in a forthcoming publication. Furthermore, the obtained HCF contains input delays, which will result in a dynamic extension, thus further analysis is required to obtain a form without delays as this would enable the application of quasi-static state feedback. Finally, the system class (15) is lacking in-domain couplings, as these were recently addressed in [13] for the MIMO case without boundary dynamics, the next logical step is to combine the two approaches.

### v3

@@ IV-D5 Algorithm
The matrix $\bar{}\mathbf{G}(\sigma,{\mathrm{d}_{t}})=\mathbf{L}(\sigma,{\mathrm{d}_{t}})\mathbf{G}(\sigma,{\mathrm{d}_{t}})$ obtained by LABEL:alg:shift_red satisfies the requirements (23).
@@
In the while loop, for the case of unique pivot degrees, the degrees in each column are corrected separately from left to right. Therein, each correction step can introduce new shifts to the rows below its pivot element. In the case of coinciding degrees of pivot elements, a matrix-valued quasipolynomial long division (MQPLD) is performed which reduces the degrees in several columns at once but is not discussed here due to space limitations. Then the loop repeats, thus, effectively reducing the delays in each column by the (nonzero) degree differences of the pivot elements until the desired state dimension is reached. ∎
IV-E Deriving a hyperbolic MIMO controller form
Applying LABEL:alg:shift_red to $\mathbf{G}(\sigma,{\mathrm{d}_{t}})$ from (20) we obtain
$\mathbf{L}^{-1}(\sigma,{\mathrm{d}_{t}})\bm{u}(t)=\bar{}\mathbf{G}(\sigma,{\mathrm{d}_{t}})\bm{y}(t)$ (31)
with $\hat{\mathbf{G}}=\operatorname{lccm}_{{\mathrm{d}_{t}},\sigma}\bar{}\mathbf{G}(\sigma,{\mathrm{d}_{t}})$ having full rank. Thus, by means of the separation $\mathbf{\bar{G}}(\sigma,{\mathrm{d}_{t}})=\mathbf{\hat{G}}\mathbf{K}(\sigma,{\mathrm{d}_{t}})+\mathbf{\check{G}}(\sigma,{\mathrm{d}_{t}})$ where $\mathbf{K}(\sigma,{\mathrm{d}_{t}})=\operatorname{diag}({\mathrm{d}_{t}}^{\nu_{1}}\sigma^{\rho_{1}},\dotsc,{\mathrm{d}_{t}}^{\nu_{m}}\sigma^{\rho_{m}})$ with the highest derivative and prediction orders $\nu_{i}=\deg_{{\mathrm{d}_{t}}}\bar{g}_{ii}(\sigma,{\mathrm{d}_{t}})$ and $\rho_{i}=\deg^{+}_{\sigma}\bar{g}_{ii}(\sigma,{\mathrm{d}_{t}})$ we obtain the explicit FDE system
$\mathbf{K}(\sigma,{\mathrm{d}_{t}})\bm{y}(t)\!=\!\hat{\mathbf{G}}^{-1}\!\!\left(\mathbf{L}^{-1}(\sigma,{\mathrm{d}_{t}})\bm{u}(t)\!-\tilde{\mathbf{G}}(\sigma,{\mathrm{d}_{t}})\bm{y}(t)\!\right)$ (32)
@@ IV-E Deriving a hyperbolic MIMO controller form
$y_{i}^{(\nu_{i})}(t+\rho_{i})=-\sum_{j=1}^{m}\left((\tilde{a}^{\star}_{ij}y_{j})(t)-(\tilde{b}^{\star}_{ij}u_{j})(t)\right)$ (33)
for $i=\{1,\dotsc,m\}$ in the analytic setting. Herein, the operators $\tilde{a}^{\star}_{ij}$ and $\tilde{b}^{\star}_{ij}$ by construction have compact support, which is why their application yields definite integrals whose limits depend on the degrees of $\tilde{\mathbf{G}}(\sigma,{\mathrm{d}_{t}})$ and $\mathbf{L}^{-1}(\sigma,{\mathrm{d}_{t}})$ w.r.t. $\sigma$, respectively. Thus, (33) states a well-posed initial value problem if the components $y_{i}(t+\tau)$ of the flat output are given for $\tau\in[\delta_{i},\rho_{i}]$ with $\delta_{i}=\deg^{-}_{\sigma}\bar{g}_{ii}(\sigma,{\mathrm{d}_{t}})$ and the input $\bm{u}(t)$ is known for $t\in\mathbb{R}$.
@@
$\eta^{\mathrm{O}}_{i,j}(t)$ $\coloneq y_{i}^{(j-1)}(t),\;j\in\{1,\dotsc,\nu_{i}\}$ (34a)
$\eta^{\mathrm{P}}_{i}(\tau,t)$ $\coloneq y_{i}^{(\nu_{i})}(t+\delta_{i}+\tau),\;\tau\in[0,\hat{}\tau_{i}]$ (34b)
with $\hat{}\tau_{i}=\rho_{i}-\delta_{i}$ for $i\in\{1,\dotsc,m\}$ which is equivalent to the flat output on these intervals. Substituting [^5] Some expressions in $\bm{y}(t)$ may remain as e.g. the values $y_{i}^{(d)}(t+\tau)$ are not part of the HCF state $\bm{\eta}(\cdot,t)$ for $d<\nu_{i}$, cf. (12). However, these expressions can be easily eliminated by means shown in [22, Lem. 5]. the definitions (34) in (33) then allows to identify the operators $a^{\star}_{ij},b^{\star}_{ij}$ as well as weights $\bm{a}_{ij}$ in (12), yielding the HCF of system (15).
@@
The proof of Lemma 2 yields $\deg_{{\mathrm{d}_{t}}}q(\sigma,{\mathrm{d}_{t}})<0$ due to the structure of $d({\mathrm{d}_{t}})$. Then by Lemma 3, $q^{\star}(\sigma,{\mathrm{d}_{t}})$ has non-positive degree w.r.t. $\mathrm{d}_{t}$} as do all entries in $\mathbf{L}_{i}(\sigma,{\mathrm{d}_{t}})$ from LABEL:alg:shift_red. Now, as the inverse of the triangular $\mathbf{L}(\sigma,{\mathrm{d}_{t}})$ is computed by forward substitution, all $\bar{q}^{\star}_{ij}(\sigma,{\mathrm{d}_{t}})$ therein are linear combinations of the $q_{ij}^{\star}(\sigma,{\mathrm{d}_{t}})$, hence, $\deg_{{\mathrm{d}_{t}}}\bar{q}^{\star}_{ij}(\sigma,{\mathrm{d}_{t}})\leq 0$. Therefore, the application of the $\tilde{b}^{\star}_{ij}$ yields (distributed) shifts but no derivatives. ∎
@@ V Reduction of the benchmark system
Consider again the system from Section II-C with ${l_{1}}=\pi$, ${l_{2}}=10$ and M=1{M}=1 substituted into (10) yielding
$\mathbf{G}(\sigma,{\mathrm{d}_{t}})=\frac{1}{4}\begin{pmatrix}a_{1}({\mathrm{d}_{t}})\sigma^{\pi}+a_{2}({\mathrm{d}_{t}})\sigma^{-\pi}&-\sigma^{\pi}-\sigma^{-\pi}\\ b_{1}({\mathrm{d}_{t}})\sigma^{10}+b_{2}({\mathrm{d}_{t}})\sigma^{-10}&\sigma^{10}+\sigma^{-10}\\ \end{pmatrix}$
with $a_{1}({\mathrm{d}_{t}})=b_{1}({\mathrm{d}_{t}})={\mathrm{d}_{t}}^{2}+2{\mathrm{d}_{t}}$ and $a_{2}({\mathrm{d}_{t}})=b_{2}({\mathrm{d}_{t}})={\mathrm{d}_{t}}^{2}-2{\mathrm{d}_{t}}$ for which $\hat{}\mathbf{G}=\operatorname{lccm}_{\sigma,{\mathrm{d}_{t}}}\mathbf{G}(\sigma,{\mathrm{d}_{t}})=\operatorname{lccm}_{{\mathrm{d}_{t}},\sigma}\mathbf{G}(\sigma,{\mathrm{d}_{t}})$ is singular as shown in (11) which is why the degree conditions (24) from Section IV-D1 fail; e.g. $\deg_{\sigma}g_{21}(\sigma,{\mathrm{d}_{t}})=20>\deg_{\sigma}g_{11}(\sigma,{\mathrm{d}_{t}})=2\pi$.
Continuing with Section IV-D2, no superfluous delays are present, rendering the delay triangularisation and homogenisation obsolete. Furthermore, the row degrees w.r.t. $\sigma$ are already ordered and the columns do not need to be reordered as their entries satisfy the degree condition (24b).
Therefore, by means of Lemma 2 the quotient
$\begin{aligned} q_{21}&(\sigma,{\mathrm{d}_{t}})\!=\!\sigma^{10-\pi}\!+\!\frac{2-{\mathrm{d}_{t}}}{{\mathrm{d}_{t}}+2}\sigma^{10-3\pi}\!-\!\frac{{\mathrm{d}_{t}}+2}{{\mathrm{d}_{t}}-2}\sigma^{-10+3\pi}\!+\!\sigma^{-10+\pi}\end{aligned}$
is obtained, where one iteration was needed for each case in (28). Next, Lemma 3 yields the correction term $p({\mathrm{d}_{t}})=\frac{16e^{-20+6\pi}}{\left({\mathrm{d}_{t}}^{2}-4\right)}$ such that $q_{21}^{\star}(\sigma,{\mathrm{d}_{t}})\in\mathcal{R}_{\mathrm{e}}$. Using the latter in (26) then yields the reduced element
$\begin{aligned} \bar{g}_{21}(\sigma,{\mathrm{d}_{t}})=\!-\frac{16{\mathrm{d}_{t}}}{\left({\mathrm{d}_{t}}-2\right)e^{20-6\pi}}\sigma^{\pi}\!+\frac{{\mathrm{d}_{t}}\left({\mathrm{d}_{t}}^{2}+4{\mathrm{d}_{t}}+4\right)}{{\mathrm{d}_{t}}-2}\sigma^{-10+4\pi}\\ \!+\frac{{\mathrm{d}_{t}}\left({\mathrm{d}_{t}}^{2}-4{\mathrm{d}_{t}}+4\right)}{{\mathrm{d}_{t}}+2}\sigma^{10-4\pi}\!-\frac{16{\mathrm{d}_{t}}}{\left({\mathrm{d}_{t}}+2\right)e^{20-6\pi}}\sigma^{-\pi}.\end{aligned}$
with $\deg_{\sigma}\bar{g}_{21}(\sigma,{\mathrm{d}_{t}})=\deg_{\sigma}g_{11}(\sigma,{\mathrm{d}_{t}})$. As the requirements for the second column are already satisfied, this concludes the reduction and we obtain $\mathbf{\hat{G}}=\operatorname{lccm}_{{\mathrm{d}_{t}},\sigma}\mathbf{\bar{G}}(\sigma,{\mathrm{d}_{t}})=\operatorname{diag}(1,-2)$.  [^6] The complete computation of the HCF for the example with annotated steps can be found at https://github.com/umit-iace/iace-pub-2026-ecc.
@@ VI Concluding remarks
This contribution establishes a HCF for MIMO systems with dynamic boundaries and proposes a flatness based algorithm to compute said form for a specific class of systems. While the algebraic approach taken to obtain this result is limited to linear systems, it enables a systematic procedure which lends well to an implementation using computer algebra. However, the insight gained from an analytic interpretation of the algebraic procedure allows for the technique to be extended to nonlinear boundary systems, which will be discussed in a forthcoming publication. Furthermore, the obtained HCF contains input delays, which will result in a dynamic extension, thus further analysis is required to obtain a form without delays as this would enable the application of quasi-static state feedback. Finally, the system class (15) is lacking in-domain couplings. As this class was recently addressed in [23] without boundary dynamics, the next logical step is to combine the two approaches.